Papenfuss Family Inn is a bed and breakfast establishment in a converted 100-year-old mansion. The Inn's guests appreciate its gourmet breakfasts and individually decorated rooms. The Inn's overhead budget for the most recent month appears below:

Activity level 62 guests
Variable overhead costs:
Supplies $ 83.70
Laundry 409.20
Fixed overhead costs:
Utilities 540.00
Salaries and wages 5,140.00
Depreciation 2,650.00
Total overhead cost $ 8,822.90


The Inn's variable overhead costs are driven by the number of guests.
What would be the total budgeted overhead cost for a month if the activity level is 52guests?

$53,317.40

$8,822.90

$8,743.40

$7,842.24
Answer #1

Total Variable cost per guest = (83.70+409.20)/62

= 7.95 per guest

Total fixed cost = 540+5140+2650 = 8330

Total budgeted overhead cost

= Variable (7.95*52) + Fixed 8330

= 8743.40

Option C